Evaluate the following. Tel(x- 5y + 2) ds S: 2 = 13 - 6x + 8y, OS X 51,0 Sys 3 X
Answer #1

- I 03751 02423 1: n=0 Y=0 dy 3 I D) (4-5y+z) ds Z=13-6*+By 1 Scu-5y + 13- 6x+8y)* ds- site in the or andy z=13-67 toy ( 3y -